# Deeply virtual $\phi$-meson production near threshold

### v1

@@ II Kinematics
[…]
@@
At HERA, exclusive $\phi$-production has been measured at high energy, up to $W\sim 130$ GeV and $Q^{2}\lesssim 40$ GeV2 Adloff et al. (2000); Chekanov et al. (2005); Aaron et al. (2010), primarily motivated by the physics of diffraction at $x_{B}\ll 1$. Such vector meson electroproduction data have been utilized for modeling GPDs at LO Goloskokov and Kroll (2007); Meskauskas and Müller (2014) and, more recently, as inputs for global analyses of GPDs at NLO in deeply virtual Compton scattering (DVCS) and DVMP Lautenschlager et al. (2013); Čuić et al. (2023). Turning to the low energy side, the LEPS collaboration Mibe et al. (2005) has measured the photoproduction ($Q=0$) cross section near the threshold $W_{th}\leq W\leq 2.31$ GeV ($1.57\leq E_{\gamma}\leq 2.37\,{\rm GeV}$ in terms of the photon energy in the target (proton) rest frame). Subsequently, the CLAS collaboration Santoro et al. (2008) has measured the electroproduction cross section in the range $W_{th}<W<3$ GeV and $3.81.4<Q^{2}<$ 3.8 GeV2. Currently, there are several experimental proposals at JLab to measure electroproduction near the threshold at higher $Q^{2}$, and the effort can be continued in the future with the EIC.
@@ III.1 General consideration
Finally, near the threshold, the relative velocity between the produced $\phi$ and the recoiling proton is small and can be nonrelativistic (zero at the threshold by definition). This means that there could be final state interactions (FSIs) which potentially break factorization Hatta and Strikman (2021). From a perturbative QCD perspective, the standard collinear factorization theorem Collins et al. (1997) assumes that the outgoing proton with momentum is collinear to the incoming proton
@@
As soon as $\xi\gtrsim 1/3$, this becomes a stronger constraint than $Q^{2}\gg|t|$ and further shrinks the region of applicability toward $t\sim t_{min}$ and $\xi\lesssim 0.6$. However, whether this effect is numerically important in practice is largely unknown. Ultimately, what matters is the interaction strength between the produced meson and the recoiling proton, which depends on the meson species. In this regard, we refer to a recent dedicated study Kim et al. (2024) where it has been concluded that, near the threshold, the $\phi+p$ FSI cross section is orders of magnitude smaller than the production cross section. While Ref. Kim et al. (2024) concerns photoproduction, we expect the same conclusion holds in electroproduction since the final state is the same. We therefore neglect final-state interactions and stick to (22), although more investigations are certainly desirable.
@@ III.2 Conformal partial wave expansion
where ${\cal H}$ and ${\cal E}$ are the DVMP amplitudes which contain the GPDs $H^{q,g}$ and $E^{q,g}$, respectively. $q$ is the label for quark flavors. Throughout this paper, we consider four active flavors $n_{f}=4$, or $q=u,d,s,c$. ${\cal H}$ and ${\cal E}$ have the following factorized structure
@@
where $e_{s}=-\frac{1}{3}$ is the electric charge (in units of $|e|$) of the $s$-quark. The notations $N_{c}=3$, $C_{F}=\frac{N_{c}^{2}-1}{2N_{c}}=\frac{4}{3}$ are standard. $\varphi(u)$ and $f_{\phi}$ are the $\phi$-meson distribution amplitude (DA) and the decay constant, respectively. $T^{a}$ are the hard scattering amplitudes, also referred to as coefficient functions. The symbols $\otimes$ represent convolutions in $-1<x<1$ and $10<u<$ 1. For simplicity, we have chosen the factorization scales of GPD and DA to be the same and equal to the renormalization scale $\mu^{2}$.
@@
and similarly for $E_{j}^{q,g}(\xi,t)$. The notation $H^{q(+)}(x,\xi,t)=H^{q}(x,\xi,t)-H^{q}(-x,\xi,t)$ denotes the C-even part. The gluon GPD is normalized such that, in the forward limit, $H^{g}(x,0,0)=xG(x)$, where $G(x)$ is the unpolarized gluon PDF. These moments reduce to the standard Mellin moments of PDFs in the $\xi\to 0$ limit.222 Note that the Gegenbauer polyanomials in (30), (31) are defined for any $-\infty<x/\xi<\infty$ which is outside of the Efremov-Radyushkin-Brodsky-Lepage (ERBL) region $-1<x/\xi<1$ where they form a set of orthogonal polynomials. Compare with (38) where the latter condition is enforced by $\theta(\xi-|x|)$. The first moments $j=1$ are the quark and gluon GFFs (1)
@@ III.3 Threshold approximation
deferring the detailed numerical analysis to the next section. If we compare only the $A_{q,g}$ terms and assume a similar $t$-dependence for the quark and gluon GFFs, clearly the gluon exchange dominates in ${\cal H}$. The NLO correction for this term ${\cal O}(\alpha_{s}^{2})$ is sizable, about 60-70% of the leading order result. We also notice that the contribution from the $s$-quark exchange is comparable to that from $u,d$-quarks despite $A_{s}\ll A_{u+d}$ because the latter only enters at NLO. The relative minus sign suggests a significant cancellation between the two contributions. However, the inclusion of the $D_{q,g}$ terms may drastically change this expectation. Since $\xi^{2}\sim{\cal O}(1)$ and $D_{q,g}$ are supposedly negative, there can be a significant cancellation in $A_{q,g}+\xi^{2}D_{q,g}$ if $D_{q,g}$ are order unity. However, this cancellation may not be effective in the $s$-quark sector because $A_{s}\ll 1$. Therefore, if $D_{s}\sim{\cal O}(1)$, the $s$-quark contribution could become dominant Hatta and Strikman (2021). $D_{q,g}$ play an even more important role in ${\cal E}$ since there are indications that $B_{q,g}$ are very small Teryaev (1999); Hagler et al. (2008). As is clear from (26), when $\xi\sim{\cal O}(1)$, the ${\cal E}$ term is no longer negligible relative to the ${\cal H}$ term.
@@ III.4 Accuracy of the threshold approximation
Let us investigate for which values of $\xi$ the threshold approximation (III.3), (III.3) can be a decent approximation to the full DVMP amplitude. We shall focus on the GPD $H$ and the associated amplitude ${\cal H}$, but the discussion for $E$ and ${\cal E}$ is entirely analogous. In the asymptotic limit $\mu^{2}\to\infty$ (corresponding to $Q^{2}\to\infty$) where $H_{g}(x,\xi)\propto\theta(\xi-|x|)(1-\frac{x^{2}}{\xi^{2}})^{2}$, the $j=1$ conformal moment by definition accounts for 100% of the LO amplitude, irrespective of the value of $\xi$. (In the Mellin moment expansion, the $j=1$ moment accounts for 80% of the amplitude in the limit $\xi\to 1$ Hatta and Strikman (2021); Guo et al. (2021).) If $\mu^{2}$ is finite this is not the case, however. Indeed, the inverse transforms of (30) and (31) are given by the formal series
$H^{q(+)}(x,\xi,t)$`="$\sum_{j=1}^{\infty}\frac{2^{j+1}\Gamma(5/2+j)}{\xi^{j+1}\Gamma(3/2)\Gamma(3+j)}\left(1-\frac{x^{2}}{\xi^{2}}\right)H_{j}^{q}(\xi,t)C_{j}^{3/2}(x/\xi)\Theta(\xi-|x|),$ (38)
$H^{g}(x,\xi,t)$`="$\sum_{j=1}^{\infty}\frac{3}{\xi^{j}}\frac{2^{j+1}\Gamma(5/2+j)}{\Gamma(3/2)\Gamma(4+j)}\left(1-\frac{x^{2}}{\xi^{2}}\right)^{2}H_{j}^{g}(\xi,t)C_{j-1}^{5/2}(x/\xi)\Theta(\xi-|x|),$ (39)
where the symbol ``="``"``="` ` = " indicates that the summation is generally divergent, unless the GPDs have support only in the ERBL region $-\xi<x<\xi$, which is not the case for finite $\mu^{2}$. However, one may as well expand the GPDs in terms of the basis $C_{j}^{3/2}(x)$, $C_{j}^{5/2}(x)$ defined over their correct support region $-1<x<1$ Belitsky and Radyushkin (2005). The modified moments are given by
@@
On general grounds, one expects that convergent orthogonal series are dominated by low moments because higher moments involve stronger oscillating functions which destructively interfere with more slowly varying functions like GPDs. It then remains to be seen whether the convolution with the coefficient functions $T^{a}(x,\xi)$ preserves this expectation. In order to quantitatively test this, we use the Goloskokov-Kroll (GK) model Goloskokov and Kroll (2007, 2008) for $H^{q,g}$ with PDF parameters fitted to the PDF sets Alekhin et al. (2017, 2018). [^3] This model has been used for the numerical analysis in Braun et al. (2022). The GK model was originally designed for the small-$x_{B}$, small-$t$ region. Assuming an exponential $t$-dependence in GPDs, the model successfully describes the HERA data on meson production cross sections. However, it has not been tested in the region of our interest, namely, $\xi\sim{\cal O}(1)$ where $|t|\geq|t_{min}|=\frac{4\xi^{2}M^{2}}{1-\xi^{2}}$ becomes large and one starts to see the power-law behavior in $t$ Kroll et al. (2013). Given the potentially strong model dependence at large-$t$, in the following we use the GK model evaluated at an unphysical point $H^{q,g}(x,\xi,t=0)$ (see also Guo et al. (2024)) which can be constructed relatively unambiguously using the double distribution technique Radyushkin (1999). We will comment on the finite-$t$ effect later. Fig. 3 shows $H^{u,d,s,g}(x,\xi,t=0,\mu^{2})$ at a representative point $\xi=0.5$ and $\mu=2$ GeV.
@@
The contours $\gamma_{j}$ can be defined as straight lines with respect to an arbitrary point $z\in\mathbb{C}$ with $\text{Im}\,z>0$ in the following way. $\gamma_{1}$ goes from 00 to $z$, $\gamma_{2}$ goes from $\xi$ to $z$ and $\gamma_{3}$ goes from $z$ to $1$. In practice, $z$ should be chosen for optimal numerical stability.
We then compare the resulting LO and NLO amplitudes ${\cal H}$ with their truncated versions ${\cal H}_{\rm trunc}$ (III.3) separately in the three channels ${\cal H}={\cal H}^{s}+{\cal H}^{ps}+{\cal H}^{g}$ corresponding to the three terms in (III.3). (The superscript $ps$ stands for ‘pure singlet’ which enters only at NLO, see the middle diagram of Fig. 2.) This is plotted in Fig. 4 in terms of the relative errors
@@
It can be seen that the truncation error in the $s$-quark contribution is the largest, at best 20% around $\xi\approx 0.5$. The error does not decrease by increasing $\xi$, indicating that higher moments are important. Still, the situation is much better than in the Mellin moment expansion where the truncation error can be much larger Hatta and Strikman (2021). For the gluon and pure-singlet contributions, the threshold approximation is very good, with errors being below 10% starting already at $\xi\approx 0.35$. Interestingly, the approximation is the best for the total amplitude ${\cal H}$ for $\xi>0.5$, which may be a special feature of $\phi$-production. This is due to a significant cancellation between the $s$-quark and pure singlet contributions both in the truncated and full amplitudes, see the discussion below (37) and Fig. 5 where $\mathcal{H}^{s}_{\rm NLO}\approx-\mathcal{H}^{ps}_{\rm NLO}$ in the entire $\xi$ range. As a result, $\mathcal{H}$ is almost entirely dominated by the gluon contribution $\mathcal{H}\approx\mathcal{H}^{g}$. We expect that the threshold approximation gets even better as $\mu^{2}\sim Q^{2}$ is increased, since the GPDs approach their asymptotic limits.
An important question is whether the above conclusion, obtained at an unphysical point $t=0$, remains qualitatively valid in the phenomenologically relevant region $t\approx t_{min}$. Empirically, it is known that the gluon GPD and form factors fall faster with increasing $|t|$ than the quark counterparts Lukashin et al. (2001); Chekanov et al. (2005); Aaron et al. (2008). One can then imagine a scenario where the $\sum_{q}A_{q}$ term in (III.3) cancels the $A_{g}$ term (instead of the $A_{s}$ term as mentioned above) at large-$t$. If this rather accidental cancellation occurs, $R^{\rm tot}$ can become very large $|R^{\rm tot}|\gg 1$ due to a small denominator $|{\cal H}_{\rm trunc}|$ even though the threshold approximation works well for individual contributions $|R^{s,ps,g}|\ll 1$. In fact, this is what happens if one evaluates ${\cal H}$ using $H(x,\xi,t=t_{min}(\xi))$ and the default parameter set of the GK model Goloskokov and Kroll (2008). We however think that this is an artifact of the GK model which predicts a too rapid falloff of the gluon GPD and GFFs at large-$t$ as compared to the quark ones. A more reasonable estimate of the $t$-dependence may come from lattice QCD Hackett et al. (2024), where it was found that $\sum_{q}A_{q}(t)$ and $A_{g}(t)$ have quite similar $t$-dependencies. A simple numerical check using (37) and the parametrization of $A_{q,g}(t)$ obtained in Hackett et al. (2024) shows that the $A_{g}$ term safely dominates over the $\sum_{q}A_{q}$ term in (III.3) in the phenomenologically relevant region $|t|<2$ GeV2. Therefore, for the moment we exclude the above scenario, although further investigations are certainly needed.
@@
Overall, we conclude that errors due to the threshold approximation are at worst $20\%$ and can be as good as $10\%$ or less for moderately large $\xi\gtrsim 0.4$. Considering also the discussion around (25), it appears that the range $0.60.4<\xi<$ 0.6 is the best region to focus on.
@@ IV Numerical analysis
In this section we present our numerical results on the differential cross section $\frac{d\sigma_{L}}{dt}$ (26), where ${\cal H},{\cal E}$ are given by (III.3), (III.3). The inputs for the NLO computation are as follows. The one-loop running coupling is
@@
At first sight, this is surprising because superficially the formula (26) suggests that, at fixed $W$,444 Note that, in the usual Bjorken limit $Q^{2}\to\infty$ at fixed $x_{B}$, $\frac{d\sigma_{L}}{dt}\sim Q^{-6}$ up to logarithms in $Q$, although this has tension with the HERA data at $x_{B}\ll 1$, see the discussion in Čuić et al. (2023).
@@ V Experimental Projections for the EIC and JLab
[…]
@@ V.1 EIC
[…]
The momentum spectra of all four final-state particles in reconstructed near-threshold events are shown in Fig. 9. The scattered electron and the kaons from the $\phi$ decay are required to be within the central detector acceptance $|\eta|<3.5$. In near-threshold events, the scattered proton is most often measured in the B0 detector system, which spans approximately $5.94.6<\eta<$ 5.9. However, the azimuthal angle acceptance of the B0 is not 100% for all polar angles. To compensate for the missing azimuthal acceptance, the assumed detection efficiency of the B0 detectors is set to 70%. Finally, at the smallest scattering angles, the proton can be detected in the Roman Pots (RP) or Off-momentum Detectors (OMD), which reside inside the beam vacuum. Since these detectors rely on particles passing through the accelerator magnet optics along certain trajectories, they do not have acceptance for all momenta. To crudely simulate this effect, we include a longitudinal momentum dependent acceptance of $10.65<p_{z}/p_{\mathrm{beam}}<$ 1 and $0.60.3<p_{z}/E_{\mathrm{beam}}<$ 0.6. There will also be a dependence of the acceptance on $p_{T}$, which we make no attempt to capture here. We also note that the above numbers hold for the 18x275 magnet lattice, but the acceptance will likely change in the 5x41 magnet lattice. A second detector at the EIC will have a still-different lattice and may make use of a second focus point in the accelerator magnets to significantly enhance the RP/OMD acceptance. The present assumption of 95% efficiency in the RP/OMD is optimistic. However, for the range of $W$ that we focus on, namely $2.82.4<W<$ 2.8, less than 10% of scattered protons are produced at $\eta>5.9$. We therefore expect our results to hold regardless of the details of the RP/OMD acceptance.
@@
Central $(|\eta|<3.5)$ 95% $\frac{\Delta p}{p}=\begin{cases}1\%,&|\eta|<2.5\\ 2.5\%,&2.5<|\eta|<3.5\\ \end{cases}$ $\sigma_{\theta,\phi}=3\,\text{mrad\quad}$
@@
Figure 9: Momentum (Left) and pseudorapidity (Right) distributions for particles in events with $41.96<W<$ 4 GeV and $Q^{2}>2$ GeV2 where all four final-state particles are reconstructed.
@@
The current projection for the EIC is that the integrated luminosity per year will be 5 fb-1 for the 5x41 beam energy configuration. We assume in our projections a total integrated luminosity of 10 fb-1, corresponding to approximately two years of 5x41 running. The 10x100 beam energy configuration offers a factor of 10 higher instantaneous luminosity, making it an attractive option. However, for the 10x100 beam energy, in the near-threshold kinematics the produced kaons almost exclusively fall in the pseudorapidity region $4.63.5<\eta<$ 4.6. In ePIC this pseudorapidity range is occupied primarily by the beampipe and is not instrumented with tracking and paricle identification. A second EIC detector with forward tracking and PID acceptance out to $\eta=4$ or greater could recover a significant fraction of these events and perhaps do a more precise measurement at the 10x100 beam energy.
@@
Figure 11: Projected results at the EIC for the cross section $\gamma^{*}_{L}+p\rightarrow\phi+p$ differential in $|t|$ and $Q^{2}$ in the near-threshold regime of $2.82.4<W<$ 2.8. The data points are fixed to the central value of the theoretical curve. Theoretical uncertainties are determined via a scale variation in the same way as Fig. 6. The assumed integrated luminosity is 10 fb-1 at the 5x41 beam energy.
The projected sensitivity to $d\sigma_{L}/d|t|$ at the EIC is shown in Fig. 11. For simplicity, the NLO theory curves are generated at fixed values of $W$ and $Q^{2}$ corresponding to the center of the bin, i.e. $W=2.6$ GeV and $Q^{2}=3.5$ GeV2 for the lowest $Q^{2}$ bin. The reconstructed fully exclusive event yield is around 1500 in 10 fb-1 for $2.82.4<W<$ 2.8 GeV and $Q^{2}>3$ GeV2. In all bins, the statistical uncertainty is larger than the assumed 10% systematic uncertainty. We present our results only in the range $2.82.4<W<$ 2.8 GeV because this range provides the best discriminating power for $D_{s}$. At lower $W$, the event statistics are too small to perform a reliable measurement of the $|t|$-distribution. At larger $W$, despite the somewhat higher measured event statistics, the sensitivity to $D_{s}$ is diminished. Besides, as explained before, our calculation is not reliable for $W>3$ GeV. Therefore we predict that $2.82.4<W<$ 2.8 GeV will be an optimal range for this measurement if the B0 detector meets the performance presented in Table 1 and the acceptance of the forward particle identification is not significantly reduced compared to the nominal $\eta<3.5$ boundary.

### v3

@@ II Kinematics
$\sigma^{\gamma^{*}p}=\int^{t_{max}}_{t_{min}}dt\frac{d\sigma^{\gamma^{*}p}}{dt}=\sigma_{T}(W,Q)+\varepsilon\sigma_{L}(W,Q)=\left(\frac{1}{R(W,Q)}+\varepsilon\right)\sigma_{L}(W,Q),$ (16)
@@
At HERA, exclusive $\phi$-production has been measured at high energy, up to $W\sim 130$ GeV and $Q^{2}\lesssim 40$ GeV2 Adloff et al. (2000); Chekanov et al. (2005); Aaron et al. (2010), primarily motivated by the physics of diffraction at $x_{B}\ll 1$. Such vector meson electroproduction data have been utilized for modeling GPDs at LO Goloskokov and Kroll (2007); Meskauskas and Müller (2014) and, more recently, as inputs for global analyses of GPDs at NLO in deeply virtual Compton scattering (DVCS) and DVMP Lautenschlager et al. (2013); Čuić et al. (2023). Turning to the low energy side, the LEPS collaboration Mibe et al. (2005) has measured the photoproduction ($Q=0$) cross section near the threshold $W_{th}\leq W\leq 2.31$ GeV ($1.57\leq E_{\gamma}\leq 2.37\,{\rm GeV}$ in terms of the photon energy in the target (proton) rest frame). Subsequently, the CLAS collaboration Santoro et al. (2008) has measured the electroproduction cross section in the range $W_{th}<W<3$ GeV and $1.4<Q^{2}<3.8$ GeV2. Currently, there are several experimental proposals at JLab to measure electroproduction near the threshold at higher $Q^{2}$, and the effort can be continued in the future with the EIC.
@@ III.1 General consideration
Finally, near the threshold, the relative velocity between the produced $\phi$ and the recoiling proton is small and can be nonrelativistic (zero at the threshold by definition). This means that there could be final state interactions (FSIs) which potentially break factorization Hatta and Strikman (2021). From a perturbative QCD perspective, the standard collinear factorization theorem Collins et al. (1997) assumes that the outgoing proton is collinear to the incoming proton
@@
As soon as $\xi\gtrsim 1/3$, this becomes a stronger constraint than $Q^{2}\gg|t|$ and further shrinks the region of applicability toward $t\sim t_{min}$ and $\xi\lesssim 0.6$. However, whether this effect is numerically important in practice is largely unknown. Ultimately, what matters is the interaction strength between the produced meson and the recoiling proton, which depends on the meson species. In this regard, we refer to a recent dedicated study Kim et al. (2021, 2024) where it has been concluded that, near the threshold, the $\phi+p$ FSI cross section is orders of magnitude smaller than the production cross section. While Ref. Kim et al. (2024) concerns photoproduction, we expect the same conclusion holds in electroproduction since the final state is the same. We therefore neglect final-state interactions and stick to (22), although more investigations are certainly desirable.
@@ III.2 Conformal partial wave expansion
[…]
@@
where $e_{s}=-\frac{1}{3}$ is the electric charge (in units of $|e|$) of the $s$-quark The notations $N_{c}=3$, $C_{F}=\frac{N_{c}^{2}-1}{2N_{c}}=\frac{4}{3}$ are standard. $\varphi(u)$ and $f_{\phi}$ are the $\phi$-meson distribution amplitude (DA) and the decay constant, respectively. $T^{a}$ are the hard scattering amplitudes, also referred to as coefficient functions. The symbols $\otimes$ represent convolutions in 1<x<1-1<x<1 and $0<u<1$. For simplicity, we have chosen the factorization scales of GPD and DA to be the same and equal to the renormalization scale $\mu^{2}$.
@@
and similarly for $E_{j}^{q,g}(\xi,t)$. The notation $H^{q(+)}(x,\xi,t)=H^{q}(x,\xi,t)-H^{q}(-x,\xi,t)$ denotes the C-even part. The gluon GPD is normalized such that, in the forward limit, $H^{g}(x,0,0)=xG(x)$, where $G(x)$ is the unpolarized gluon PDF. These moments reduce to the standard Mellin moments of PDFs in the $\xi\to 0$ limit. [^2] Note that the Gegenbauer polyanomials in (30), (31) are defined for any $\infty<x/\xi<\infty$ which is outside of the Efremov-Radyushkin-Brodsky-Lepage (ERBL) region $1<x/\xi<1$ where they form a set of orthogonal polynomials. Compare with (38) where the latter condition is enforced by $\theta(\xi-|x|)$. The first moments $j=1$ are the quark and gluon GFFs (1)
@@ III.3 Threshold approximation
[…]
@@ III.4 Accuracy of the threshold approximation
$H^{q(+)}(x,\xi,t)$ $``="\sum_{j=1}^{\infty}\frac{2^{j+1}\Gamma(5/2+j)}{\xi^{j+1}\Gamma(3/2)\Gamma(3+j)}\left(1-\frac{x^{2}}{\xi^{2}}\right)H_{j}^{q}(\xi,t)C_{j}^{3/2}(x/\xi)\Theta(\xi-|x|),$ (38)
$H^{g}(x,\xi,t)$ $``="\sum_{j=1}^{\infty}\frac{3}{\xi^{j}}\frac{2^{j+1}\Gamma(5/2+j)}{\Gamma(3/2)\Gamma(4+j)}\left(1-\frac{x^{2}}{\xi^{2}}\right)^{2}H_{j}^{g}(\xi,t)C_{j-1}^{5/2}(x/\xi)\Theta(\xi-|x|),$ (39)
where the symbol =``=" indicates that the summation is generally divergent, unless the GPDs have support only in the ERBL region $\xi<x<\xi$, which is not the case for finite $\mu^{2}$. However, one may as well expand the GPDs in terms of the basis $C_{j}^{3/2}(x)$, $C_{j}^{5/2}(x)$ defined over their correct support region 1<x<1-1<x<1 Belitsky and Radyushkin (2005). The modified moments are given by
$\mathbb{H}^{q}_{j}(\xi,t)=\frac{\Gamma(3/2)\Gamma(j+1)}{2^{j}\Gamma(j+3/2)}\frac{1}{2}\int_{-1}^{1}dx\,C_{j}^{3/2}(x)H^{q(+)}(x,\xi,t),$ (40)
@@
On general grounds, one expects that convergent orthogonal series are dominated by low moments because higher moments involve stronger oscillating functions which destructively interfere with more slowly varying functions like GPDs. It then remains to be seen whether the convolution with the coefficient functions $T^{a}(x,\xi)$ preserves this expectation. In order to quantitatively test this, we use the Goloskokov-Kroll (GK) model Goloskokov and Kroll (2007, 2008) for $H^{q,g}$ with PDF parameters fitted to the PDF sets Alekhin et al. (2017, 2018). [^3] This model has been used for the numerical analysis in Braun et al. (2022). The GK model was originally designed for the $x_{B}, small$-$t$ region. Assuming an exponential $t$-dependence in GPDs, the model successfully describes the HERA data on meson production cross sections. However, it has not been tested in the region of our interest, namely, $\xi\sim{\cal O}(1)$ where $|t|\geq|t_{min}|=\frac{4\xi^{2}M^{2}}{1-\xi^{2}}$ becomes large and one starts to see the power-law behavior in $t$ Kroll et al. (2013). Given the potentially strong model dependence at large-$t$, in the following we use the GK model evaluated at an unphysical point $H^{q,g}(x,\xi,t=0)$ (see also Guo et al. (2024)) which can be constructed relatively unambiguously using the double distribution technique Radyushkin (1999). We will comment on the finite-$t$ effect later. Fig. 3 shows $H^{u,d,s,g}(x,\xi,t=0,\mu^{2})$ at a representative point $\xi=0.5$ and $\mu=2$ GeV.
@@
$\int_{0}^{1}dx\,T(x,\xi-i\epsilon)H(x,\xi)=\int_{\gamma_{1}}dx\,T(x,\xi)H_{2}(x,\xi)+\int_{\gamma_{2}}dx\,T(x,\xi)(H_{1}(x,\xi)-H_{2}(x,\xi))+\int_{\gamma_{3}}dx\,T(x,\xi)H_{1}(x,\xi).$ (43)
The contours $\gamma_{j}$ can be defined as straight lines with respect to an arbitrary point $z\in\mathbb{C}$ with $\text{Im}\,z>0$ in the following way. $\gamma_{1}$ goes from 0 to $z$, $\gamma_{2}$ goes from $\xi$ to $z$ and $\gamma_{3}$ goes from $z$ to $1$. In practice, $z$ should be chosen for optimal numerical stability.
@@
[…]
@@
Overall, we conclude that errors due to the threshold approximation are at worst $20\%$ and can be as good as $10\%$ or less for moderately large $\xi\gtrsim 0.4$. Considering also the discussion around (25), it appears that the range $0.4<\xi<0.6$ is the best region to focus on.
@@ IV Numerical analysis
[…]
@@
At first sight, this is surprising because superficially the formula (26) suggests that, at fixed WW, [^4] Note that, in the usual Bjorken limit $Q^{2}\to\infty$ at fixed $x_{B}$, $\frac{d\sigma_{L}}{dt}\sim Q^{-6}$ up to logarithms in $Q$, although this has tension with the HERA data at $x_{B}\ll 1$, see the discussion in Čuić et al. (2023).
@@ V Experimental Projections for the EIC and JLab
In practice, an actual experiment will measure the $e+p$ cross section within a certain kinematic range and then extract the longitudinal cross section using the formula (16). While one may use a model for $R$ as in our simulations, $R$ can be experimentally determined if the spin-density matrix elements (SDMEs) of the $\phi$ are measured. [^6] RR may also be measured via the Rosenbluth separation technique, wherein cross sections are measured at the same kinematics ($x_{B},Q^{2},t$) but different values of $\varepsilon$. In practice, $\varepsilon$ is typically varied by altering the beam energy. Measurements of this kind are well suited to high-luminosity spectrometer experiments where the kinematics are known precisely and point-to-point uncertainties are small. A measurement of $R$ in exclusive $\phi$ electroproduction using the Rosenbluth method would provide a valuable cross-check of the SDME method. However, no Rosenbluth separation of exclusive $\phi$ production has been performed to date. Assuming $s$-channel helicity conservation (SCHC), the value of $R$ is given as
@@ V.1 EIC
The ePIC experiment will be operational at the beginning of EIC running and located at the 6 o’clock interaction point. ePIC offers excellent acceptance and capability for measuring the momentum and species of long-lived particles such as pions, kaons, and protons. To make projections for measurements of exclusive $\phi$ electroproduction we will assume acceptances, resolutions, and particle identification abilities similar to what should be achieved by ePIC, summarized in Table 1 ePIC Collaboration (2025). However, it is worth noting that a second EIC detector, coming online after ePIC and the first phase of EIC running, could offer improved capabilities for these measurements. The configuration of the EIC which collides 5 GeV electrons with 41 GeV protons (5x41) [^7] In the following, the notation AAxBB will refer to collisions of electrons with momentum $A$ with protons of momentum $B$. offers the best experimental access to the near-threshold region because the charged kaons produced in the decay of the $\phi$ enter the central detector region of $3.5<\eta<3.5$. These $\phi$ decay kaons predominantly populate the forward region of the detector, where ePIC is equipped with a dual-radiator Ring-Imaging Cherenkov (dRICH) detector capable of efficiently separating pions from kaons up to approximately 50 GeV Chatterjee (2024). For the purposes of this study, we assume that the dRICH and time-of-flight system can provide perfect identification of kaons from $\phi$ decay, which have typical momenta of 5 GeV.
The momentum spectra of all four final-state particles in reconstructed near-threshold events are shown in Fig. 9. The scattered electron and the kaons from the $\phi$ decay are required to be within the central detector acceptance $|\eta|<3.5$. In near-threshold events, the scattered proton is most often measured in the B0 detector system, which spans approximately $4.6<\eta<5.9$. However, the azimuthal angle acceptance of the B0 is not 100% for all polar angles. To compensate for the missing azimuthal acceptance, the assumed detection efficiency of the B0 detectors is set to 70%. Finally, at the smallest scattering angles, the proton can be detected in the Roman Pots (RP) or Off-momentum Detectors (OMD), which reside inside the beam vacuum. Since these detectors rely on particles passing through the accelerator magnet optics along certain trajectories, they do not have acceptance for all momenta. To crudely simulate this effect, we include a longitudinal momentum dependent acceptance of $0.65<p_{z}/p_{\mathrm{beam}}<1$ and $0.3<p_{z}/E_{\mathrm{beam}}<0.6$. There will also be a dependence of the acceptance on $p_{T}$, which we make no attempt to capture here. We also note that the above numbers hold for the 18x275 magnet lattice, but the acceptance will likely change in the 5x41 magnet lattice. A second detector at the EIC will have a still-different lattice and may make use of a second focus point in the accelerator magnets to significantly enhance the RP/OMD acceptance. The present assumption of 95% efficiency in the RP/OMD is optimistic. However, for the range of $W$ that we focus on, namely $2.4<W<2.8$, less than 10% of scattered protons are produced at $\eta>5.9$. We therefore expect our results to hold regardless of the details of the RP/OMD acceptance.
@@
[…]
@@
Figure 9: Momentum (Left) and pseudorapidity (Right) distributions for particles in events with $1.96<W<4$ GeV and $Q^{2}>2$ GeV2 where all four final-state particles are reconstructed.
@@
The current projection for the EIC is that the integrated luminosity per year will be 5 fb-1 for the 5x41 beam energy configuration. We assume in our projections a total integrated luminosity of 10 fb-1, corresponding to approximately two years of 5x41 running. The 10x100 beam energy configuration offers a factor of 10 higher instantaneous luminosity, making it an attractive option. However, for the 10x100 beam energy, in the near-threshold kinematics the produced kaons almost exclusively fall in the pseudorapidity region $3.5<\eta<4.6$. In ePIC this pseudorapidity range is occupied primarily by the beampipe and is not instrumented with tracking and paricle identification. A second EIC detector with forward tracking and PID acceptance out to $\eta=4$ or greater could recover a significant fraction of these events and perhaps do a more precise measurement at the 10x100 beam energy.
@@
Figure 11: Projected results at the EIC for the cross section $\gamma^{*}_{L}+p\rightarrow\phi+p$ differential in $|t|$ and $Q^{2}$ in the near-threshold regime of $2.4<W<2.8$. The data points are fixed to the central value of the theoretical curve. Theoretical uncertainties are determined via a scale variation in the same way as Fig. 6. The assumed integrated luminosity is 10 fb-1 at the 5x41 beam energy.
The projected sensitivity to $d\sigma_{L}/d|t|$ at the EIC is shown in Fig. 11. For simplicity, the NLO theory curves are generated at fixed values of $W$ and $Q^{2}$ corresponding to the center of the bin, i.e. $W=2.6$ GeV and $Q^{2}=3.5$ GeV2 for the lowest $Q^{2}$ bin. The reconstructed fully exclusive event yield is around 1500 in 10 fb-1 for $2.4<W<2.8$ GeV and $Q^{2}>3$ GeV2. In all bins, the statistical uncertainty is larger than the assumed 10% systematic uncertainty. We present our results only in the range $2.4<W<2.8$ GeV because this range provides the best discriminating power for $D_{s}$. At lower $W$, the event statistics are too small to perform a reliable measurement of the $|t|$-distribution. At larger $W$, despite the somewhat higher measured event statistics, the sensitivity to $D_{s}$ is diminished. Besides, as explained before, our calculation is not reliable for $W>3$ GeV. Therefore we predict that $2.4<W<2.8$ GeV will be an optimal range for this measurement if the B0 detector meets the performance presented in Table 1 and the acceptance of the forward particle identification is not significantly reduced compared to the nominal $\eta<3.5$ boundary.